In Amy’s backyard, there are also chickens and rabbits. There are 5 more chickens than rabbits. The total number of legs is 130. How many rabbits and how many chickens?
Answer by Boreal(15235)

x rabbits with 4x legs
x+5 chickens with 2x+10 legs
4x+2x+10=130
6x+10=130
6x=120
x=20 rabbits, 80 legs
x+5=25 chickens, 50 legs
130 legs